OpenStudy (anonymous):

Which circles have a circumference of 20pie units? Choose all answers that are correct. A. a circle with a diameter of 40 units B. a circle with a radius of 10 units C. a circle with a diameter of 20 units D. a circle with a radius of 20 units

OpenStudy (anonymous):

mmmm pie...

OpenStudy (anonymous):

o.o

OpenStudy (anonymous):

circumference is \(2\pi r\)

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

A. a circle with a diameter of 40 units if the diameter is \(40\) then the radius is \(20\) (half the diameter) so not that one, since its circumference would be \(2\times 20\times \pi=40\pi\)

OpenStudy (anonymous):

try the next one

OpenStudy (anonymous):

only A?

OpenStudy (anonymous):

@starlord6200

OpenStudy (anonymous):

It might also be b jelly

OpenStudy (anonymous):

I'm jelly your peanut

OpenStudy (anonymous):

Yea it's b 2

OpenStudy (anonymous):

Yea i chose B to